OpenStudy (anonymous):

Let f(x) = x – 2 and g(x) = 3x + 4. What is f(g(x))

OpenStudy (anonymous):

f(g(x)) means you plug in g(x) for x in f(x): f(g(x)) = g(x) - 2 = 3x + 4 - 2 = 3x + 2.

OpenStudy (anonymous):

thank you